OpenStudy (anonymous):

solve the following equation: 5x+√x+7 -13=0

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 5x+ \sqrt{x+7 }-13=0 }\) \(\LARGE\color{blue}{ \bf \sqrt{x+7 }=13-5x }\) \(\LARGE\color{blue}{ \bf x+7 =169-130x+25x^{2} }\) \(\LARGE\color{blue}{ \bf 0 =162-131x+25x^{2} }\) solve it from here....

OpenStudy (anonymous):

thank you…got it